User:Brian P. Josey/Notebook/2009/11/24
|Project name||Main project page|
Previous entry Next entry
So the PCR from yesterday worked. The primers that we used were F834-dig and R 2008. This differs from last weeks primers, in that the reverse primer was changed. As with before, the only differences between the nine different samples was the concentration of Mg2+ cations. The concentrations ranged from 1.0 mM to 5.0 mM in 0.5 mM increments. They are on the gel in increaseing concentration running from left to right. As you can see 3.5 mM and 5.0 mM did not work. All the others however did.
I set up the digestion of pBR322 today, to get it ready for tomorrow. To do this we mixed 20 μL of the DNA, pBR322, 2.5 μL of buffer # 4 and 2.5 μL of the EarI enzyme. We then warmed this at 37°C for an hour and put it in the -20°C freezer. If anyone needs it, it will be on the black rack and labled with pBR322. The detailed set up is:
For Wednesday 50mL of the TAE buffer in jar and 0.4mg of agarose for the gel and big comb. (purple thing)